Can you help me solve (((3x^4-28x^2+9=0)))
Answer by stanbon(75887)

(3x^2-1)(x^2-9)=0
3x^2-1=0 or x^2-9=0
x^2=(1/3) or x=3 or x=-3
x=sqrt3/3 or x=-sqrt3/3 or x=3 or x=-3
Cheers,
Stan H.